OpenStudy (aravindg):

let f(x)=3 sinx and g(x)=cos^2 x find (f+g)(pie/3)

OpenStudy (anonymous):

What, again...?

OpenStudy (anonymous):

didnt someone just answer it :s?

OpenStudy (aravindg):

giv the answer i didnt get it

OpenStudy (anonymous):

3sinx+cos^x is the equation okay now sub pie/3 in it

OpenStudy (aravindg):

thn??

OpenStudy (aravindg):

wat is final answer? so i can verify

OpenStudy (anonymous):

2.84 i think

OpenStudy (aravindg):

thx a lot

OpenStudy (aravindg):

let f(x)=x^2 and g(x)=x^3 define fg.........

OpenStudy (anonymous):

(x^3)(x^2)=x^5

OpenStudy (aravindg):

is that the answer???

OpenStudy (anonymous):

yes